TIFFGT(1) TIFFGT(1) NAME tiffgt−displayanimagestoredinaTIFFfile(SiliconGraphicsversion) SYNOPSIS tiffgt[options]input.tif... DESCRIPTION tiffgtdisplaysoneormoreimagesstoredusingtheTagImageFileFormat,Revision6.0.Each imageisplacedinafixedsizewindowthattheusermustpositiononthedisplay(unlessconfiguredotherwisethroughXdefaults).Ifthedisplayhasfewerthan24bitplanes,oriftheimage doesnotwarrantfullcolor,thenRGBcolorvaluesaremappedtotheclosestvaluesthatexistin thecolormap(thisisdoneusingthergbiroutinefoundinthegraphicsutilitylibrary−lgutil.) tiffgtcorrectlyhandlesfileswithanyofthefollowingcharacteristics: BitsPerSample1,2,4,8,16 SamplesPerPixel1,3,4(the4thsampleisignored) PhotometricInterpretation0(min-is-white),1(min-is-black),2(RGB),3(palette),6(YCbCr) PlanarConfiguration1(contiguous),2(separate) Orientation1(top-left),4(bottom-left) Datamaybeorganizedasstripsortilesandmaybecompressedwithanyofthecompressionalgorithmssupportedbythelibtiff(3)library. Forpaletteimages(PhotomatricInterpretation=3),tiffgtinspectsthecolormapvaluesandassumes either16-bitor8-bitvaluesaccordingtothemaximumvalue.Thatis,ifnocolormapentry greaterthan255isfound,tiffgtassumesthecolormaphasonly8-bitvalues;otherwiseitassumes 16-bitvalues.Thisinspectionisdonetohandleoldimageswrittenbyprevious(incorrect)ver-sionsoflibtiff. tiffgtcanbeusedtodisplaymultipleimagesone-at-a-time.Theleftmousebuttonswitchesthe displaytothefirstimageinthenextfileinthelistoffilesspecifiedonthecommandline.The rightmousebuttonswitchestothefirstimageinthepreviousfileinthelist.Themiddlemouse buttoncausesthefirstimageinthefirstfilespecifiedonthecommandlinetobedisplayed.In additionthefollowingkeyboardcommandsarerecognized: bUseaPhotometricIntepretationofMinIsBlackindisplayingthecurrentimage. lUseaFillOrderoflsb-to-msbindecodingthecurrentimage. mUseaFillOrderofmsb-tolmsbindecodingthecurrentimage. cUseacolormapvisualtodisplaythecurrentimage. rUseatruecolor(24-bitRGB)visualtodisplaythecurrentimage. wUseaPhotometricIntepretationofMinIsWhiteindisplayingthecurrentimage. WToggle(enable/disable)displayofwarningmessagesfromtheTIFFlibrarywhendecoding images. EToggle(enable/disable)displayoferrormessagesfromtheTIFFlibrarywhendecoding images. zResetallparameterstotheirdefaultsettings(FillOrder,PhotometricInterpretation,handlingofwarningsanderrors). PageUp Displaythepreviousimageinthecurrentfileorthelastimageinthepreviousfile. PageDown Displaythenextimageinthecurrentfileorthefirstimageinthenextfile. HomeDisplaythefirstimageinthecurrentfile. October15,1995 1 TIFFGT(1) TIFFGT(1) EndDisplaythelastimageinthecurrentfile(unimplemented). OPTIONS −cForceimagedisplayinacolormapwindow. −dSpecifyanimagetodisplaybydirectorynumber.Bydefaultthefirstimageinthefileis displayed.Directoriesarenumberedstartingatzero. −eEnablereportingoferrormessagesfromtheTIFFlibrary.Bydefaulttiffgtsilentlyignores imagesthatcannotberead. −fForcetiffgttorunasaforegroundprocess.Bydefaulttiffgtwillplaceitselfinthebackgroundonceithasopenedtherequestedimagefile. −lForcethepresumedbitorderingtobeLSBtoMSB. −mForcethepresumedbitorderingtobeMSBtoLSB. −oSpecifyanimagetodisplaybydirectoryoffset.Bydefaultthefirstimageinthefileis displayed.DirectoriesoffsetsmaybespecifiedusingC-stylesyntax;i.e.aleading‘‘0x’’for hexadecimalandaleading‘‘0’’foroctal. −pOverridethevalueofthePhotometricInterpretationtag;theparametermaybeoneof: miniswhite,minisblack,rgb,palette,mask,separated,ycbcr,andcielab. −rForceimagedisplayinafullcolorwindow. −sStoponthefirstreaderror.Bydefaultallerrorsintheinputdataareignoredandtiffgt doesit’sbesttodisplayasmuchofanimageaspossible. −wEnablereportingofwarningmessagesfromtheTIFFlibrary.Bydefaulttiffgtignores warningmessagesgeneratedwhenreadinganimage. −vPlaceinformationinthetitlebardescribingwhattypeofwindow(fullcolororcolormap) isbeingused,thenameoftheinputfile,andthedirectoryindexoftheimage(ifnonzero).Bydefault,thewindowtypeisnotshowninthetitlebar. BUGS Imageswiderandtallerthanthedisplayaresilentlytruncatedtoavoidcrashingoldversionsof thewindowmanager. SEEALSO tiffdump(1),tiffinfo(1),tiffcp(1),libtiff(3) October15,1995 2